A cylinder of Ne gas has a volume of 9.76L. It contains 61.2g Ne at 23.0⁰C. What is the pressure in the cylinder?

Please show how you solved this
Answer #1

Molar mass of Ne = 20.18 g/mol

mass(Ne)= 61.2 g

use:

number of mol of Ne,

n = mass of Ne/molar mass of Ne

=(61.2 g)/(20.18 g/mol)

= 3.033 mol

Given:

V = 9.76 L

n = 3.0327 mol

T = 23.0 oC

= (23.0+273) K

= 296 K

use:

P * V = n*R*T

P * 9.76 L = 3.0327 mol* 0.08206 atm.L/mol.K * 296 K

P = 7.55 atm

Answer: 7.55 atm